# This is the extended version of the conference version at FoSSaCS 2022 [BFPZ22] [a] [a] [a]

### v2

@@ 1. Introduction
In recent years, scientific and technological advancement in computer science and engineering led to an ever increasing influence of computer systems on our everyday lives. A lot of decisions which were historically done by humans are now in the hands of intelligent systems. At the same time, these systems grow more and more complex, and thus, harder to understand. This poses a huge challenge in the development of reliable and trustworthy systems. Therefore, an important task of computer science today is the broad development of comprehensive and versatile ways to understand modern software and cyber physical systems.
@@
As we like to to have probability-raising inherent in the MDP, we require (C1) under every scheduler. Thus, the cause-effect relation holds for every resolution of the non-deterministic choices. We consider two natural ways to interpret condition (C1): On one hand, the probability-raising property can be required locally for each element of the cause. This results in a strict property which requires that after each execution leading to the cause the probability of effect has been raised. Such causes are called strict probability-raising (SPR) causes in our framework. This interpretation is especially suited when the task is to identify system states that have to be avoided for lowering the effect probability. On the other hand, one might want to treat the cause globally as a unit in (C1) leading to the notion of global probability-raising (GPR) cause. This way, the causal relation can also be formulated between properties instead of considering individual elements or executions to be causal. Considering the cause as a whole is also better suited when further constraints are imposed on the candidates for cause sets. E.g. if the set of non-terminal states of the given MDP is partitioned into sets of states $S_{i}$ under control of different agents $i$, $1\leq i\leq k$ and the task is to identify which agent’s decisions might cause the effect, only the subsets of $S_{1},\ldots,S_{k}$ are candidates for causes. Furthermore, global causes are more appropriate when causes are used for monitoring purposes under partial observability constraints as then the cause candidates are sets of indistinguishable states.
Even with the distinction between strict and global probability-raising causality, different causes might still vary substantially regarding how well they predict the effect. Within Markov decision processes this intuitively coincides with how well the executions exhibiting the cause cover executions showing the effect. However, solely focusing on broader coverage might also lead to more trivial causal relations. In order to take this trade off into account, we take inspiration from measures for binary classifiers used in statistical analysis (see, e.g., [Pow11]) to introduce quality measures for causes. These allow us to compare different causes and to look for optimal causes: The recall captures the probability that the effect is indeed preceded by the cause. The coverage-ratio quantifies the fraction of the probability that cause and effect are observed and the probability that the effect but not the cause is observed. The f-score, a widely used quality measure for binary classifiers, is the harmonic mean of recall and precision, where the precision is the probability that the cause is followed by the effect. Finally, we address the computation of arbitrary quality measures as long as they can be represented as algebraic functions.
@@
The survey article [BDF+21] introduces notions of global probability-raising causes for Markov chains, where causes and effects can be path properties. [BDF+21]’s notion of reachability causes in Markov chains directly corresponds to our notion GPR causes, the only difference being that [BDF+21] deals with a relaxed minimality condition and requires that the cause set is reachable without visiting an effect state before. The latter is inherent in our approach as we suppose that all states are reachable and the effect states are terminal. On the other hand if we restrict [BDF+21]’s notion of global PR-cause to $\omega$-regular effects and co-safety causes, this corresponds to our notion of a co-safety GPR cause with the exception of the minimality condition. The same can be said about the correspondence of local PR-causes from [BDF+21] and co-safety SPR causes.
@@ 2.1. Markov decision processes
A Markov decision process (MDP) is a tuple $\mathcal{M}=(S,\mathit{Act},P,\mathsf{init})$ where $S$ is a finite set of states, $\mathit{Act}$ a finite set of actions, $\mathsf{init}\in S$ a unique initial state and $P:S\times\mathit{Act}\times S\to[0,1]$ the transition probability function such that $\sum_{t\in S}P(s,\alpha,t)\in\{0,1\}$ for all states $s\in S$ and actions $\alpha\in\mathit{Act}$.
For $\alpha\in\mathit{Act}$ and $T\subseteq S$, $P(s,\alpha,T)$ is a shortform notation for $\sum_{t\in t}P(s,\alpha,t)$. An action $\alpha$ is enabled in state $s\in S$ if $\sum_{t\in S}P(s,\alpha,t)=1$. We define $\mathit{Act}(s)=\{\alpha\in\mathit{Act}\mid\alpha\text{ is enabled in }s\}$. A state $t$ is terminal if $\mathit{Act}(t)=\emptyset$. A Markov chain (MC) is a special case of MDP where $\mathit{Act}$ is a singleton (we then write $P(s,t)$ rather than $P(s,\alpha,t)$). A path in an MDP $\mathcal{M}$ is a (finite or infinite) alternating sequence $\pi=s_{0}\,\alpha_{0}\,s_{1}\,\alpha_{1}\,s_{2}\dots\in(S\times\mathit{Act})^{*}\times S\cup(S\times\mathit{Act})^{\omega}$ such that $P(s_{i},\alpha_{i},s_{i+1})>0$ for all indices $i$. A path is called maximal if it is infinite or finite and ends in a terminal state. If $\pi$ is a finite path in $\mathcal{M}$ then $\mathit{last}(\pi)$ denotes the last state of $\pi$. That is, if $\pi=s_{0}\,\alpha_{0}\ldots\alpha_{n-1}\,s_{n}$ then $\mathit{last}(\pi)=s_{n}$.
@@
A scheduler $\mathfrak{S}$ of $\mathcal{M}$ induces a (possibly infinite) Markov chain. We write $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}$ for the standard probability measure on measurable sets of maximal paths in the Markov chain induced by $\mathfrak{S}$ with initial state $s$. If $\varphi$ is a measurable set of maximal paths, then $\mathrm{Pr}^{\max}_{\mathcal{M},s}(\varphi)$ and $\mathrm{Pr}^{\min}_{\mathcal{M},s}(\varphi)$ denote the supremum resp. infimum of the probabilities for $\varphi$ under all schedulers. We use the abbreviation $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},\mathsf{init}}$ and notations $\mathrm{Pr}^{\max}_{\mathcal{M}}$ and $\mathrm{Pr}^{\min}_{\mathcal{M}}$ for extremal probabilities. Analogous notations will be used for expectations. So, if $f$ is a random variable, then, e.g., $\mathrm{E}^{\mathfrak{S}}_{\mathcal{M}}(f)$ denotes the expectation of $f$ under $\mathfrak{S}$ and $\mathrm{E}^{\max}_{\mathcal{M}}(f)$ its supremum over all schedulers. We also use conditional probabilities in MDPs cf. [BKKM14, Mär20]. For two measurable sets of maximal paths $\varphi$ and $\psi$ we have
$\mathrm{Pr}^{\max}_{\mathcal{M},s}(\varphi\mid\psi)=\max_{\mathfrak{S}}\ \mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}(\varphi\mid\psi)=\max_{\mathfrak{S}}\ \frac{\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}(\varphi\wedge\psi)}{\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}(\psi)}$
and $\mathrm{Pr}^{\min}_{\mathcal{M},s}(\varphi\mid\psi)$ likewise. For reachability objectives maximal (and minimal) conditional probabilities are computable in polynomial time [BKKM14].
@@ 2.3. MEC-quotient
      The $\tau$-action stands for the deterministic transition to the fresh state $\bot$, i.e.: $P^{\prime}(s_{\mathcal{E}_{i}},\tau,\bot)=1.$ Suppose now that $u\in U_{i}$ and $\alpha\in\mathit{Act}_{\mathcal{M}}(u)\setminus\mathcal{E}_{i}$. Then, we set $P^{\prime}(s_{\mathcal{E}_{i}},\alpha,s^{\prime})=P(u,\alpha,\iota^{-1}(s^{\prime}))$ for all $s^{\prime}\in S^{\prime}$.
@@
The original MDP and its MEC-quotient have been found to be connected by the following lemma (see also [dA97, dA99]).
@@
For the formal definition of $\mathfrak{T}$, we use the following notation. For simplicity, let us assume that $\mathsf{init}\notin U_{1}\cup\ldots\cup U_{k}$. This yields $\mathsf{init}=\mathsf{init}^{\prime}$. Given a finite path $\pi=s_{0}\,\alpha_{0}\,s_{1}\,\alpha_{1}\ldots\alpha_{m-1}\,s_{m}$ in $\mathcal{M}$ with $s_{0}=\mathsf{init}$, let $\pi_{\mathcal{N}}$ the path in $\mathcal{N}$ resulting from by replacing each maximal path fragment $s_{h}\alpha_{h}\ldots\alpha_{j-1}s_{j}$ consisting of actions inside an $\mathcal{E}_{i}$ with state $s_{\mathcal{E}_{i}}$. (Here, maximality means if $h>0$ then $\alpha_{h-1}\notin\mathcal{E}_{i}$ and if $j<m$ then $\alpha_{j+1}\notin\mathcal{E}_{i}$.) Furthermore, let $p_{\pi}^{\mathfrak{S}}$ denote the probability for $\mathfrak{S}$ to generate the path $\pi$ when starting in the first state of $\pi$.
@@ 2.4. Automata and $\omega$-regular languages
In order to represent an $\omega$-regular language, we use deterministic Rabin automata (DRA). A DRA is a tuple $\mathcal{A}=(Q,\Sigma,q_{0},\delta,\mathrm{Acc})$ where $Q$ is a finite set of states, $\Sigma$ an alphabet, $q_{0}$ the initial state, $\delta:Q\times\Sigma\to Q$ the transition function and $\mathrm{Acc}\subseteq 2^{Q}\times 2^{Q}$ the acceptance set. The run of $\mathcal{A}$ on a word $w=w_{0}w_{1}\dots\in\Sigma^{\omega}$ is the sequence $\tau=q_{0}q_{1}\ldots$ of states such that $\delta(q_{i},w_{i})=q_{i+1}$ for all $i$. It is accepting if there exists a pair $(L,K)\in\mathrm{Acc}$ such that $L$ is only visited finitely often and $K$ is visited infinitely often by $\tau$. The language $\mathcal{L}(\mathcal{A})$ is the set of all words $w\in\Sigma^{\omega}$ on which the run of $\mathcal{A}$ is accepting.
@@
The regular language of minimal good prefixes of a co-safety $\mathcal{L}$ which uniquely determines $\mathcal{L}$ can be represented by a deterministic finite automaton (DFA). A DFA is a tuple $\mathcal{A}=(Q,\Sigma,q_{0},\delta,\mathrm{Acc})$ where $Q$ is a finite set of states, $\Sigma$ an alphabet, $q_{0}$ the initial state, $\delta:Q\times\Sigma\to Q$ the transition function and $\mathrm{Acc}\subseteq Q$ the acceptance set. The run of $\mathcal{A}$ on a finite word $w=w_{0}w_{1}\dots w_{n}$ is the sequence $\tau=q_{0}q_{1}\ldots q_{n}$ of states such that $\delta(q_{i},w_{i})=q_{i+1}$ for all $i$. It is accepting if $q_{n}\in\mathrm{Acc}$. The language $\mathcal{L}(\mathcal{A})$ is the set of all words $w\in\Sigma^{*}$ on which the run of $\mathcal{A}$ is accepting.
@@ 3. Strict and global probability-raising causes
Our contribution starts by providing novel formal definitions for cause-effect relations in MDPs which rely on the probability-raising (PR) principle $P(E\mid C)>P(C)$ stated by (C1) and temporal priority of causes as stated by (C2) in the introduction. Here, we focus on the case where both causes and effects are state properties, i.e., sets of states.
@@ 3.1. Examples and simple properties of probability-raising causes
Figure 2. A MC in which no PR cause exists
@@
Suppose, the task is to check whether taking action $\alpha$ in state $s$ raises the effect probabilities compared to never scheduling $\alpha$ in state $s$. This form of action causality was discussed in an example in [DFT20]. We argue that we can deal with this kind of causality to. For this we assume there are no cycles in $\mathcal{M}$ containing $s$. Let $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$ be copies of $\mathcal{M}$ with the following modifications: In $\mathcal{M}_{0}$, the only enabled action of state $s$ is $\alpha$, while in $\mathcal{M}_{1}$ the enabled actions of state $s$ are the elements of $\mathit{Act}_{\mathcal{M}}(s)\setminus\{\alpha\}$.
The action $\alpha$ raises the effect probability in $\mathcal{M}$ if and only if for all scheduler $\mathfrak{S}$ of $\mathcal{N}$ the copy of $s$ in $\mathcal{M}_{0}$ satisfies (SPR) in $\mathcal{N}$. This idea can be generalized to check whether scheduler classes satisfying a regular constraint have higher effect probability compared to all other schedulers. In this case, we can deal with an MDP $\mathcal{N}$ as above where $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$ are defined as the synchronous product of deterministic finite automata and $\mathcal{M}$.
@@
$\mathrm{Pr}_{\mathcal{N}}^{\mathfrak{S}}(\lozenge\mathsf{eff}\mid\lozenge s_{0})=1/4>1/8=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\lozenge\mathsf{eff})$,

### v6

@@ 1. Introduction
In recent years, scientific and technological advancement in computer science and engineering led to an ever increasing influence of computer systems on our everyday lives. A lot of decisions which were historically done by humans are now in the hands of intelligent systems. At the same time, these systems grow more and more complex, and thus, harder to understand. This poses a huge challenge in the development of reliable and trustworthy systems. Therefore, an important task of computer science today is the development of comprehensive and versatile ways to understand modern software and cyber physical systems.
@@
As the probability-raising is inherent in the MDP, we require (C1) under every scheduler. Thus, the cause-effect relation holds for every resolution of non-deterministic choices. We consider two natural ways to interpret condition (C1): On one hand, the probability-raising property can be required locally for each element of the cause. This results in a strict property which requires that after each execution showing the cause the probability of effect has been raised. Such causes are strict probability-raising (SPR) causes in our framework. This interpretation is especially suited when the task is to identify system states that have to be avoided for lowering the effect probability. On the other hand, we can treat the cause globally as a unit in (C1) leading to the notion of global probability-raising (GPR) cause. This way, the causal relation can also be formulated between properties without considering individual elements or executions to be causal. Considering the cause as a whole is also better suited when further constraints are imposed on the candidates for cause sets. E.g. if the set of non-terminal states of the given MDP is partitioned into sets of states $S_{i}$ under control of different agents $i$, $1\leq i\leq k$ and the task is to identify which agent’s decisions might cause the effect, only the subsets of $S_{1},\ldots,S_{k}$ are candidates for causes. Furthermore, GPR causes are more appropriate when causes are used for monitoring purposes under partial observability as then the cause candidates are sets of indistinguishable states.
Despite the distinction between strict and global probability-raising causality, different causes can still vary substantially in how well they predict the effect. Within Markov decision processes this intuitively coincides with how well the executions exhibiting the cause cover executions showing the effect. However, solely focusing on broader coverage might lead to more trivial causal relations. In order to take this trade off into account, we take inspiration from measures for binary classifiers used in statistical analysis (see, e.g., [Pow11]) to introduce quality measures for causes. These allow us to compare different causes and to look for optimal causes: The recall captures the probability that the effect is indeed preceded by the cause. The coverage-ratio is the fraction of the probability that cause and effect are observed and the probability of the effect without a cause. The f-score, a widely used quality measure for binary classifiers, is the harmonic mean of recall and precision, where the precision is the probability that the cause is followed by the effect. Finally, we address the computation of quality measures which can be represented as algebraic functions.
@@
The survey article [BDF+21] introduces notions of global probability-raising causes for Markov chains, where causes and effects can be path properties. [BDF+21]’s notion of reachability causes in Markov chains directly corresponds to our notion of PR causes, the only difference being that [BDF+21] deals with a relaxed minimality condition and requires that the cause set is reachable without visiting an effect state before. The latter is inherent in our approach as we suppose that all states are reachable and the effect states are terminal. On the other hand if we restrict [BDF+21]’s notion of global PR-cause to $\omega$-regular effects and co-safety causes, this corresponds to our notion of a co-safety GPR cause with the exception of the minimality condition. The same can be said about the correspondence of local PR-causes from [BDF+21] and co-safety SPR causes.
@@ 2.1. Markov decision processes
A Markov decision process (MDP) constitutes a tuple $\mathcal{M}=(S,\mathit{Act},P,\mathsf{init})$ where $S$ is a finite set of states, $\mathit{Act}$ a finite set of actions, $\mathsf{init}\in S$ a unique initial state and $P:S\times\mathit{Act}\times S\to[0,1]$ the transition probability function such that $\sum_{t\in S}P(s,\alpha,t)\in\{0,1\}$ for all states $s\in S$ and actions $\alpha\in\mathit{Act}$.
For $\alpha\in\mathit{Act}$ and $T\subseteq S$, $P(s,\alpha,T)$ is a shortform notation for $\sum_{t\in t}P(s,\alpha,t)$. An action $\alpha$ is enabled in state $s\in S$ if $\sum_{t\in S}P(s,\alpha,t)=1$. Then, $\mathit{Act}(s)=\{\alpha\in\mathit{Act}\mid\alpha\text{ is enabled in }s\}$. A state $t$ is terminal if $\mathit{Act}(t)=\emptyset$. A Markov chain (MC) is a special case of MDP where $\mathit{Act}$ is a singleton (we then write $P(s,t)$ rather than $P(s,\alpha,t)$). A path in an MDP $\mathcal{M}$ is a (finite or infinite) alternating sequence $\pi=s_{0}\,\alpha_{0}\,s_{1}\,\alpha_{1}\,s_{2}\dots\in(S\times\mathit{Act})^{*}\times S\cup(S\times\mathit{Act})^{\omega}$ such that $P(s_{i},\alpha_{i},s_{i+1})>0$ for all indices $i$. A path is called maximal if it is infinite or finite and ends in a terminal state. If $\pi$ is a finite path in $\mathcal{M}$ then $\mathit{last}(\pi)$ denotes the last state of $\pi$. That is, if $\pi=s_{0}\,\alpha_{0}\ldots\alpha_{n-1}\,s_{n}$ then $\mathit{last}(\pi)=s_{n}$.
@@
A scheduler $\mathfrak{S}$ of $\mathcal{M}$ induces a (possibly infinite) Markov chain. We write $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}$ for the standard probability measure on measurable sets of maximal paths in the Markov chain induced by $\mathfrak{S}$ with initial state $s$. If $\varphi$ is a measurable set of maximal paths, then $\mathrm{Pr}^{\max}_{\mathcal{M},s}(\varphi)$ and $\mathrm{Pr}^{\min}_{\mathcal{M},s}(\varphi)$ denote the supremum resp. infimum of the probabilities for $\varphi$ under all schedulers. We use the abbreviation $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},\mathsf{init}}$ and notations $\mathrm{Pr}^{\max}_{\mathcal{M}}$ and $\mathrm{Pr}^{\min}_{\mathcal{M}}$ for extremal probabilities. Analogous notations will be used for expectations. So, if $f$ is a random variable, then, e.g., $\mathrm{E}^{\mathfrak{S}}_{\mathcal{M}}(f)$ denotes the expectation of $f$ under $\mathfrak{S}$ and $\mathrm{E}^{\max}_{\mathcal{M}}(f)$ its supremum over all schedulers.
We also use conditional probabilities in MDPs cf. [BKKM14, Mär20]. For two measurable sets of maximal paths $\varphi$ and $\psi$ we have
$\mathrm{Pr}^{\max}_{\mathcal{M},s}(\varphi\mid\psi)=\max_{\mathfrak{S}}\ \mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}(\varphi\mid\psi)=\max_{\mathfrak{S}}\ \frac{\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}(\varphi\wedge\psi)}{\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}(\psi)},$
where $\mathfrak{S}$ ranges over all scheduler for which $\mathrm{Pr}^{\mathfrak{S}}(\psi)>0$. We define $\mathrm{Pr}^{\min}_{\mathcal{M},s}(\varphi\mid\psi)$ likewise. If both $\varphi$ and $\psi$ are reachability properties then maximal conditional probabilities are computable in polynomial time [BKKM14]. The proposed algorithm for maximal conditional probabilities $\mathrm{Pr}^{\max}_{\mathcal{M},\mathsf{init}}(\lozenge G\mid\lozenge F)$ relies on a model transformation generating a new MDP $\mathcal{N}$ that distinguishes the modes ”before $G$ and FF” (where $\mathcal{N}$ essentially behaves as $\mathcal{M}$ with additional reset transitions from end components to the initial state), ”before $G$, after FF” (where $\mathcal{N}$ maximizes the probability to reach $G$), ”before $F$, after GG” (where $\mathcal{N}$ maximizes the probability to reach $F$). Essentially the reset transitions serve to “discard” paths that never reach $G$ and $F$. For further details we refer to [BKKM14, MBKK17, Mär20].
@@ 2.3. MEC-quotient
      The $\tau$-action is the deterministic transition to the fresh state $\bot$, i.e.: $P^{\prime}(s_{\mathcal{E}_{i}},\tau,\bot)=1.$ For $u\in U_{i}$ and $\alpha\in\mathit{Act}_{\mathcal{M}}(u)\setminus\mathcal{E}_{i}$ we set $P^{\prime}(s_{\mathcal{E}_{i}},\alpha,s^{\prime})=P(u,\alpha,\iota^{-1}(s^{\prime}))$ for all $s^{\prime}\in S^{\prime}$.
@@
The original MDP and its MEC-quotient have been found to be connected by the following lemma (see also [dA97, dA99]). For the sake of completeness we present the proof for our version of the MEC-quotient.
@@
For the formal definition of $\mathfrak{T}$, we use the following notation. For simplicity, let us assume that $\mathsf{init}\notin U_{1}\cup\ldots\cup U_{k}$. This yields $\mathsf{init}=\mathsf{init}^{\prime}$. Consider a finite path $\pi=s_{0}\,\alpha_{0}\,s_{1}\,\alpha_{1}\ldots\alpha_{m-1}\,s_{m}$ in $\mathcal{M}$ with $s_{0}=\mathsf{init}$, then $\pi_{\mathcal{N}}$ is the path in $\mathcal{N}$ resulting from by replacing each maximal path fragment $s_{h}\alpha_{h}\ldots\alpha_{j-1}s_{j}$ consisting of actions inside an $\mathcal{E}_{i}$ with state $s_{\mathcal{E}_{i}}$. (Here, maximality means if $h>0$ then $\alpha_{h-1}\notin\mathcal{E}_{i}$ and if $j<m$ then $\alpha_{j+1}\notin\mathcal{E}_{i}$.) Furthermore, let $p_{\pi}^{\mathfrak{S}}$ denote the probability for $\mathfrak{S}$ to generate the path $\pi$ when starting in the first state of $\pi$.
@@ 2.4. Automata and $\omega$-regular languages
In order have a representation of an $\omega$-regular language, we use deterministic Rabin automata (DRA). A DRA constitutes a tuple $\mathcal{A}=(Q,\Sigma,q_{0},\delta,\mathrm{Acc})$ where $Q$ is a finite set of states, $\Sigma$ an alphabet, $q_{0}$ the initial state, $\delta:Q\times\Sigma\to Q$ the transition function and $\mathrm{Acc}\subseteq 2^{Q}\times 2^{Q}$ the acceptance set. The run of $\mathcal{A}$ on a word $w=w_{0}w_{1}\dots\in\Sigma^{\omega}$ is the sequence $\tau=q_{0}q_{1}\ldots$ of states such that $\delta(q_{i},w_{i})=q_{i+1}$ for all $i$. It is accepting if there exists a pair $(L,K)\in\mathrm{Acc}$ such that $L$ is only visited finitely often and $K$ is visited infinitely often by $\tau$. The language $\mathcal{L}(\mathcal{A})$ is the set of all words $w\in\Sigma^{\omega}$ on which the run of $\mathcal{A}$ is accepting.
@@
The regular language of minimal good prefixes of a co-safety $\mathcal{L}$ which uniquely determines $\mathcal{L}$ can be represented by a deterministic finite automaton (DFA). A DFA constitutes a tuple $\mathcal{A}=(Q,\Sigma,q_{0},\delta,\mathrm{Acc})$ where $Q$ is a finite set of states, $\Sigma$ an alphabet, $q_{0}$ the initial state, $\delta:Q\times\Sigma\to Q$ the transition function and $\mathrm{Acc}\subseteq Q$ the acceptance set. The run of $\mathcal{A}$ on a finite word $w=w_{0}w_{1}\dots w_{n}$ is the sequence $\tau=q_{0}q_{1}\ldots q_{n}$ of states such that $\delta(q_{i},w_{i})=q_{i+1}$ for all $i$. It is accepting if $q_{n}\in\mathrm{Acc}$. The language $\mathcal{L}(\mathcal{A})$ is the set of all words $w\in\Sigma^{*}$ on which the run of $\mathcal{A}$ is accepting.
@@ 3. Strict and global probability-raising causes
Our contribution starts by providing novel formal definitions for cause-effect relations in MDPs which rely on the probability-raising (PR) principle $P(E\mid C)>P(C)$ (C1) which states that the probability of the effect is higher after the cause. Additionally, we include temporal priority of causes (C2), stating that causes must happen before the effect. Here, we focus on the case where both causes and effects are state properties, i.e., sets of states.
@@ 3.1. Examples and simple properties of probability-raising causes
Figure 2. A MC with no PR cause
@@
Suppose, the task is to check whether taking action $\alpha$ in state $s$ raises the effect probabilities compared to never scheduling $\alpha$ in state $s$. This form of action causality was discussed in an example in [DFT20]. We argue that we can deal with this kind of causality too.
For this we assume there are no cycles in $\mathcal{M}$ containing $s$. This is a strong assumption as we do not want to force the action $\alpha$ to be taken in the first visit to $s$ or to be always taken when visiting $s$. Therefore, our framework can not handle this kind of action causality if the state in question is part of a cycle.
Let $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$ be copies of $\mathcal{M}$ with the following modifications: In $\mathcal{M}_{0}$, the only enabled action of state $s$ is $\alpha$, while in $\mathcal{M}_{1}$ the enabled actions of state $s$ are the elements of $\mathit{Act}_{\mathcal{M}}(s)\setminus\{\alpha\}$. The MDP $\mathcal{N}$ then has a fresh initial state $\mathsf{init}$ which transitions with equal probabilities $1/2$ to the copies of $s$ in $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$. The action $\alpha$ raises the effect probability in $\mathcal{M}$ if and only if for all scheduler $\mathfrak{S}$ of $\mathcal{N}$ the copy of $s$ in $\mathcal{M}_{0}$ satisfies (SPR) for the union of effect sets of both copies in $\mathcal{N}$. This idea can be generalized to check whether scheduler classes satisfying a regular constraint have higher effect probability compared to all other schedulers. In this case, we deal with an MDP $\mathcal{N}$ as above where $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$ are defined as the synchronous product of deterministic finite automata and $\mathcal{M}$.
@@
$\mathrm{Pr}_{\mathcal{N}}^{\mathfrak{S}}(\lozenge\{\mathsf{eff}_{0},\mathsf{eff}_{1}\}\mid\lozenge s_{0})=1/4>1/8=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\lozenge\{\mathsf{eff}_{0},\mathsf{eff}_{1}\}),$